OpenStudy (mysticfury):

Check My Answer? The equation y = 55x + 85 describes the amount of money a plumber charges, where y is the total amount he charges and x is the number of hours he works. The table shows the total amount of money an electrician charges. How much does each charge per hour? Number of hours worked 1 2 4 7 Total amount charged 115 180 310 505 The plumber charges $55 per hour The electrician charges $65 per hour

OpenStudy (mysticfury):

@Astrophysics @Photon336 @mathstudent55

OpenStudy (photon336):

y = 55x+85 where x = 1 hour insert x = 1 into the equation y = 55(1)+85

OpenStudy (mysticfury):

uh huh okay 140.......

OpenStudy (mysticfury):

@Photon336

OpenStudy (photon336):

hey @MysticFury that equation doesn't match the data table

OpenStudy (mysticfury):

yes i know, because the table is what the electrician charges, not the plumber... no one seems to understand that lol

OpenStudy (mysticfury):

your supposed to use the table to figure out the electricians hourly rate

OpenStudy (photon336):

plumber charges 140 we know that already now let's find the electrician's hourly rate

OpenStudy (mysticfury):

well if you did it that way it would be simple, plumber would charge 140, and electrician would charge 115, but wait that dosent add up, since 115 + 115 isnt 180

OpenStudy (photon336):

y = 55x+85 plumber y = 65x+50 electrician

OpenStudy (mysticfury):

thats what i got

OpenStudy (photon336):

I agree with you.. it's just the way I read this equation is 55 dollars par hour plus an additional 85 dollars.

OpenStudy (mysticfury):

yeah, sorry im a little grumpy but ive been doing this all day ugh